The candy store sells two types of chocolates, one for 25 cents and the other for 40 cents. If in one day she sold 100 chocolates and took in exactly 28 dollars, how many of the 25-cent chocolates did she sell?

Answer by CharlesG2(834)

The candy store sells two types of chocolates, one for 25 cents and the other for 40 cents. If in one day she sold 100 chocolates and took in exactly 28 dollars, how many of the 25-cent chocolates did she sell?
x = 25 cent chocolates, y = 40 cent chocolates
x + y = 100
0.25 * x + 0.40 * y = 28 (plug in x = 100 - y)
0.25 (100 - y) + 0.40 * y = 28
25 - 0.25 * y + 0.40 * y = 28
0.15 * y = 3
y = 20 --> 20 of the 40 cent chocolates were sold
then 80 of the 25 cent chocolates were sold